water flows at a rate of 200 liters per second up through a conical vertical pipe. the diameter at the bottom is 240 mm and at the top is 200 mm and the length is 5 m. the pressure at the bottom is 8 bar and the pressure at the top is 7.3 bar. determine.

1.- the loss of load through the pipe

2.-express it only as a function of the existing velocity head

1
Expert's answer
2021-04-08T08:51:15-0400

The loss of load through the pipe:


"\Delta F=p_1A_1-p_2A_2\\\Delta F=0.25\pi(0.24)^2(8\cdot10^5)\\-0.25\pi(0.2)^2(7.3\cdot10^5)=13257.5\ N"